a triangular traffic Island has a base half as long as its height. Find the base and the heightif the Island has an area of 64m^2
Answer by stanbon(48514)

a triangular traffic Island has a base half as long as its height. Find the base and the heightif the Island has an area of 64m^2
------------------
b = h/2
----------------
A = (1/2)b*h
----------------
64 = (1/2)(h/2)h
64 = (h/2)^2
h/2 = 8
h = 16 meters
===========
Cheers,
Stan H.